# Role of defects in atom probe analysis of sol-gel silica

### v1

@@ 3.2 Theoretical results
Nevertheless, the presence of interstitial gallium defects leads to new intra-gap electronic states (see Figure 11b), which contribute to increased optical absorption, especially at low photon energies. The simulations show that increasing the number of Ga atoms in the cubic box above 5 (i.e. at a concentration of about 2%) leads to the formation of metallic states at the Fermi level (see cyan line in Figure 11b), which makes the material conductive at odds with our experimental observations. For 4 Ga atoms in the box, our calculations indicate a significant increase in absorption from energies of 3.5 eV (see blue line in Figure 11b), even in the absence of oxygen defects, which could explain the experimentally observed features in the UV and deep-UV regions. These defect-induced states may play a role in modifying the local electronic structure without fully metallising the material, depending on the distribution and concentration of gallium defects in the host matrix. Finally, in Figure 11a we show the absorption lineshape when one or two bridging oxygen vacancies are present near or at a distance from interstitial gallium atoms, up to a concentration of 6% in Ga. We conclude that the presence of interstitial gallium atoms together with bridging oxygen vacancies in different configurations (see Figure 11a) could also explain the light absorption in the UV and deep-UV range, in agreement with our experimental results.
@@ 4 Discussion and Conclusions
The better performance of UV and deep-UV La-APT compared to green La-APT can be attributed to the higher absorption of the silica matrices in the shorter wavelength range due to defects of different nature, such as oxygen vacancies also in the presence of interstitial impurity atoms from synthesis processes.
@@
Since the optical properties of silica can be altered during nanotip fabrication by FIB-SEM milling, as already observed for Si and GaN 20, we also need to consider the effects of gallium defects at different concentrations in the silica specimen. In addition, the enormous electric field can increase the absorption by reducing the band gap of the silica in a small region near the apex where the field is very high.
@@
As far as the chemical composition is concerned, we report deviations from the nominal composition, but the measured values are close to the nominal values at a high electric field. Due to the strong local heating caused by the absorption of laser energy by the defects, the three-dimensional resolution and chemical composition are potentially limited.
To investigate the optical absorption properties of silica matrices even in the presence of defects, we have applied density functional theory with the HSE hybrid exchange-correlation functional for a variety of defects, such as interstitial and substitutional carbon, hydrogen, sodium, and gallium atoms in different concentrations and configurations. Our thorough defect analysis has highlighted four particularly relevant cases: (i) the oxygen vacancy that disrupts the Si–O–Si network by breaking a Si–Si bond; (ii) the interstitial hydrogen defect that passivates and weakens the oxygen bonds; (iii) an oxygen vacancy with a Na2 molecule nearby; (iv) interstitial and substitutional gallium atoms, which could originate from the synthesis process of the needle-shaped silica specimens for the APT analyses, also in the presence of bridging oxygen vacancies. In the latter configuration, where the gallium atoms populate the matrix with a concentration of 2%, which can be even lower if oxygen vacancies are included in the model, absorption features around 3.5-5 eV clearly emerge, indicating a signature of the defects in the experimental spectra and their role in enabling APT analyses of silica with UV and deep-UV light.
@@
This action has received funding from the European Union under grant agreement No 10104665. Part of this work was carried out at Chalmers Materials Analysis Laboratory (CMAL). The authors gratefully acknowledge the use of the HPC facilities at UniTN and FBK.
@@
The isotherm of the silica sample is shown in Figure S.1. It can be deduced from this that the material contains both micropores and mesopores. The hysteresis loop observed in the multilayer range of the nitrogen sorption measurement is typical of silica gels and porous glasses. This hysteresis is due to a difference in the relative pressure required for adsorption or desorption. This is generally caused by a capillary condensation effect that depends on the pore radius of the material 50. An interpretation of this shape is to attribute it to the network effects of pores with narrow necks and wider bodies (in the form of ink bottles), where there is blockage of the pores that remain filled until the necks are emptied during desorption 51. The micropores were found to have an average width of 6.8 Å and the mesopores an average size of 40 Å.
@@
The fits were performed locally within a $\sim$100 ns window after the Si2+ peak maximum, which was selected to avoid interference from overlapping tails of subsequent peaks. Both ToF spectra were normalised to unity at the Si2+ peak and temporally aligned such that the onset of the thermal decay ($t_{0}$) overlapped. The initial estimates and boundaries for the fitting parameters were physically motivated: $Q$ was varied around 0.1 eV based on the expectations of CSR-derived electric fields; $\Delta T_{\text{max}}$ ranged from 65 K to 250 K and $\tau$ was constrained between 0.5 and 10 ns, depending on the typical values observed in the literature 12, 13.
The fitting was performed by iterative parameter adjustment using OriginLab, aiming for a qualitative visual agreement between data and model. No automated statistical optimisation was performed; the fits are therefore considered illustrative. For the fit in Figure 4b of the main text, the extracted values of $Q_{\text{deep-UV}}=0.12$ eV and $Q_{\text{UV}}=0.06$ eV reflect the expectation that a higher electric field of 21.5 V/nm was used for UV analyses vs. 19.7 V/nm for deep-UV analyses. The values of 65K and 250 K for $T_{max}$ were determined for the UV and deep-UV analyses, respectively.
@@
Computational details of absorption spectra
Figure S.5: Effect of finite size: average of the imaginary part of the dielectric tensor, averaged in the orthogonal Cartesian directions. Amorphous silica matrix without defects represented by a cell with $N=648$ atoms corresponding to an edge of the box $L=2.14$ nm (yellow colour) compared to a cell with $N=192$ atoms corresponding to an edge of the box $L=1.42$ nm (dark green colour). These calculations are performed with a PBE exchange-correlation functional.
In Figure S.5 we show instead the imaginary part of the dielectric tensor for two different models of silicon dioxide corresponding to two different sizes of the boxes (green colour: $L=2.14$ nm, yellow colour: $L=1.42$ nm). While the details of the main peak around 10 eV are slightly different due to the different structures, the band gap of the two systems is the same. This shows us that a box size of $L=1.42$ nm is sufficient to avoid spurious interactions between the periodic images when analysing the optical properties.
Figure S.6: Effect of the DFT functional: average of the imaginary part of the dielectric tensor along the orthogonal Cartesian directions. Amorphous silica matrix using a PBE exchange-correlation functional (dark green colour) compared to a hybrid HSE functional (red colour).
Figure S.7: Real and imaginary parts of the dielectric function for the three silicon dioxide samples in Figure 7a of the main text.
Figure S.8: Refractive index of the silica sample 1 (green line), shown in Figure 7a of the main text, in comparison to experimental data (black line).
In Figure S.6 we report instead on the investigation of two different DFT functionals in the evaluation of the imaginary part of the dielectric tensor. A hybrid functional such as HSE reproduces the bandgap of amorphous silica much more accurately than the pure PBE exchange-correlation functional.

### v2

@@ 3.2 Theoretical results
Nevertheless, the presence of interstitial gallium defects leads to new intra-gap electronic states (see Figure 11d), which contribute to increased optical absorption, especially at low photon energies. The simulations show that increasing the number of Ga atoms in the cubic box above 5 (i.e. at a concentration of about 2%) leads to the formation of metallic states at the Fermi level (see cyan line in Figure 11d), which makes the material conductive at odds with our experimental observations. For 4 interstitial Ga atoms in the box, our calculations indicate a significant increase in absorption from energies of 3.5 eV (see blue line in Figure 11d), even in the absence of oxygen defects, which could explain the experimentally observed features in the UV and deep-UV regions. These defect-induced states may play a role in modifying the local electronic structure without fully metallising the material, depending on the distribution and concentration of gallium defects in the host matrix. Finally, in Figure 11e we show the absorption lineshape when one (O def.) or two (2 O def.) bridging oxygen vacancies are present near (O def. close) or at a distance (O def. far) from interstitial gallium atoms, up to a concentration of 6% in Ga. We conclude that the presence of interstitial gallium atoms together with bridging oxygen vacancies in different configurations (see Figure 11e) could also explain the light absorption in the UV and deep-UV range, in agreement with our experimental results.
@@ 4 Discussion and Conclusions
The outcome of our theoretical and computational study is that the better performance of UV and deep-UV La-APT compared to green La-APT could be due to the higher absorption of the silica matrices in the shorter wavelength range due to defects of different nature, such as oxygen vacancies also in the presence of interstitial impurity atoms from synthesis processes.
@@
Since the optical properties of silica can be altered during nanotip fabrication by FIB-SEM milling, as already observed for Si and GaN 21, we also need to consider the effects of gallium defects at different concentrations in the silica specimen. In addition, the enormous electric field can increase the absorption by reducing the band gap of the silica in a small region near the apex where the field is very high.
@@
As far as the chemical composition is concerned, we report deviations from the nominal composition, but the measured values are closer to the nominal values at a high electric field. Due to the strong local heating caused by the absorption of laser energy by the defects, the three-dimensional resolution and chemical composition are potentially limited.
To investigate the optical absorption properties of silica matrices even in the presence of defects, we have applied density functional theory with the HSE06 hybrid exchange-correlation functional for a variety of defects, such as interstitial and substitutional carbon, hydrogen, sodium, and gallium atoms in different concentrations and configurations. Our thorough defect analysis has highlighted four particularly relevant cases: (i) the oxygen vacancy that disrupts the Si–O–Si network by breaking a Si–Si bond; (ii) the interstitial hydrogen defect that passivates and weakens the oxygen bonds; (iii) an oxygen vacancy with a Na2 molecule nearby; (iv) interstitial and substitutional gallium atoms, which could originate from the synthesis process of the needle-shaped silica specimens for the APT analyses, also in the presence of bridging oxygen vacancies. In the latter configuration, where the gallium atoms populate the matrix with a concentration of 2%, which can be even lower if oxygen vacancies are included in the model, absorption features around 3.5-5 eV clearly emerge, indicating a signature of the defects in the experimental spectra and their role in enabling APT analyses of silica with UV and deep-UV light.
@@
This action has received funding from the European Union under the Mimosa grant agreement No 10104665. Part of this work was carried out at Chalmers Materials Analysis Laboratory (CMAL). The authors gratefully acknowledge the use of the HPC facilities at UniTN and FBK.
@@
The isotherm of the silica sample is shown in Figure S.1. From this it can be deduced that the material contains both micropores and mesopores. The hysteresis loop observed in the multilayer range of the nitrogen sorption measurement is typical of silica gels and porous glasses. This hysteresis is due to a difference in the relative pressure required for adsorption or desorption. This is generally caused by a capillary condensation effect that depends on the pore radius of the material 62. An interpretation of this shape is to attribute it to the network effects of pores with narrow necks and wider bodies (in the form of ink bottles), where there is blockage of the pores that remain filled until the necks are emptied during desorption 61. It was found that the micropores have an average width of 6.8 Å and the mesopores have an average size of 40 Å.
@@
Computational details of absorption spectra
Figure S.4: Effect of the finite size of the calculation supercell on the imaginary part of the dielectric tensor, averaged in the orthogonal Cartesian directions. Amorphous silica matrix without defects represented by a cell with $N=648$ atoms corresponding to an edge of the box $L=2.14$ nm (yellow colour) compared to a cell with $N=192$ atoms corresponding to an edge of the box $L=1.42$ nm (dark green colour) and to a cell with $N=72$ atoms corresponding to an edge box of $L=1.03$ nm (red colour). These calculations are performed with a PBE exchange-correlation functional.
In Figure S.4 we show instead the imaginary part of the dielectric tensor for three different models of silicon dioxide corresponding to three different sizes of the boxes (yellow colour: $L=2.14$ nm, $N$=648 atoms; green colour: $L=1.42$ nm, $N$=192 atoms; red colour: $L=1.03$ nm, $N$=72 atoms). We note that the spectrum in red is characterised by a spurious peak between 7 and 8 eV, which is not present when larger boxes are used. Conversely, while the details of the main peak around 10 eV are slightly different due to the different structures, the band gap of the two systems having $L=1.42$ nm and $L=2.14$ nm is the same. This shows us that a box size of $L=1.42$ nm is sufficient to avoid spurious interactions between the periodic images when analysing the optical properties.
Figure S.5: Effect of the DFT functional: average of the imaginary part of the dielectric tensor along the orthogonal Cartesian directions. Amorphous silica matrix using a PBE exchange-correlation functional (dark green colour) compared to a hybrid HSE06 functional (red colour).
In Figure S.5 we report instead on the investigation of two different DFT functionals in the evaluation of the imaginary part of the dielectric tensor. A hybrid functional such as HSE06 reproduces the bandgap of amorphous silica much more accurately than the pure PBE exchange-correlation functional.
@@
The fits were performed locally within a $\sim$100 ns window after the Si2+ peak maximum, which was selected to avoid interference from overlapping tails of subsequent peaks. Both ToF spectra were normalised to unity at the Si2+ peak and temporally aligned such that the onset of the thermal decay ($t_{0}$) overlapped. The initial estimates and boundaries for the fitting parameters were physically motivated: $Q$ was varied around 0.1 eV based on the expectations of CSR-derived electric fields; $\Delta T_{\text{max}}$ ranged from 65 K to 250 K and $\tau$ was constrained between 0.5 and 10 ns, depending on the typical values observed in the literature 13, 14.
The fitting was performed by iterative parameter adjustment using OriginLab, aiming for a qualitative visual agreement between data and model. No automated statistical optimisation was performed; the fits are therefore considered illustrative. For the fit in Figure 4b of the main text, the extracted values of $Q_{\text{deep-UV}}=0.12$ eV and $Q_{\text{UV}}=0.06$ eV reflect the expectation that a higher electric field of 21.5 V/nm was used for UV analyses vs. 19.7 V/nm for deep-UV analyses. The values of 65K and 250 K for $T_{\mathrm{max}}$ were determined for the UV and deep-UV analyses, respectively.
@@
Figure S.7: Real and imaginary part of the dielectric function for the three silicon dioxide samples in Figure 7a of the main text, calculated with the exchange-correlation functional HSE06.
Figure S.8: Refractive index of the silica sample 1 (green line), shown in Figure 7a of the main text, calculated with the exchange-correlation functional HSE06 in comparison to experimental data (black line).
Dielectric properties